# Electrically switchable non-relativistic Zeeman spin splittings in collinear antiferromagnets

### v1

@@ Abstract
Magnetic or electrical manipulation of electronic spin is elementary for spin-based logic, computing, and memory, where the latter is a low-power manipulation scheme. Rashba-like spin splittings stemming from spin-orbit interaction (SOI) enable electric-field manipulation of spin, but the relativistic SOI causes spin relaxations and yields dissipative transport of spin-encoded information. Recent works suggest the occurrence of electric-field switchable non-relativistic Zeeman spin splittings (NRZSSs) in collinear antiferromagnets — allowing for electrical manipulation of spin in the non-relativistic regime; yet, a theory elucidating the mechanisms for these NRZSSs and guiding the materials discovery remains missing. Here, we develop such a theory by analyzing the symmetries of spin point groups. We highlight the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms for NRZSSs that depend linearly on electric field and are electrically switchable. First-principles calculations further confirm that LiMnPO4 and NaMnP showcase such NRZSSs. Our theory provides guidelines for discovering light-element collinear antiferromagnets with electrically switchable NRZSSs, which are promising for the design of high-performance spin-based devices.
@@
Introduction — Spintronic devices (e.g., logic, computing, and memory) are built upon the manipulation of electronic spin in solids [1, 2, 3, 4, 5]. Such a manipulation can be achieved by magnetic field or electric field, where the latter offers fascinating low-power devices by reducing the “Joule heating” [6, 7, 8]. In principle, electric-field manipulation of spin is enabled by electrically switchable spin splittings [9, 10]. Because of the relativistic spin-orbit interaction (SOI), dielectrics and ferroelectrics naturally host electrically switchable Rashba-like spin splittings [11, 12, 13, 14, 15]. As a by-product, SOI causes the so-called D’yakonov-Perel’ spin relaxation as well [16, 17, 18], which yields the spin decoherence of electrons and prevents the long-range transport of the spin-encoded information [19, 18, 20, 21]. To suppress the D’yakonov-Perel’ spin relaxation, the proposal of electrically switchable spin splittings in the non-relativistic regime is of high necessity. Recent works suggest the possibility of achieving non-relativistic Zeeman spin splittings (NRZSSs) in collinear antiferromagnets that are switchable by an electric field [22, 23, 24, 25, 26, 27, 28, 29, 30]. Yet, mechanisms for such NRZSSs are elusive and selection rules for the corresponding materials discovery remain to be explored.
@@
To address the aforementioned issues, we analyze the symmetries of spin point groups (SPGs) that are associated with collinear magnets in the non-relativistic limit. This allows us to develop a theory on electrically switchable NRZSSs in collinear antiferromagnets, and reveal the magnetoelectric and piezomagnetoelectric mechanisms for this phenomenon. Our theory provides theoretical guidelines for screening collinear antiferromagnets with electrically switchable NRZSSs. This yields the first-principles predictions that LiMnPO4 and NaMnP are two candidates with linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs, respectively.
[…]
@@
For all the $[R_{s}\left|\right|R_{l}]$ symmetry operations in ferromagnetic or ferrimagnetic SPGs, $R_{s}s_{\chi}=s_{\chi}$ is always true and $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ is not necessarily enforced [36, 37]. This means that non-relativistic collinear ferromagnets or ferrimagnets enable the spontaneous NRZSSs (see e.g., Refs. [33, 34, 35]). Unlike those in ferromagnetic or ferrimagnetic SPGs, the $[R_{s}\left|\right|R_{l}]$ symmetry operations in an antiferromagnetic SPG can be partitioned into two subsets [36, 37, 34, 35], namely,
@@
and the “$\alpha$” superscript labels the symmetry reduction caused by the $E_{\alpha}$ electric field  [^3] When $E_{\alpha}$ is not compatible with the spatial $R_{l}$ operation (i.e., $R_{l}E_{\alpha}\neq E_{\alpha}$), the application of $E_{\alpha}$ break the $[R_{s}\left|\right|R_{l}]$ symmetry operation.. When $L_{-}^{\alpha}$ becomes empty, there will be no symmetry operations that enforce the $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ degeneracy — enabling the electric-field induced NRZSSs. The $E_{\alpha}$-field-induced NRZSSs are described by the effective Hamiltonian
@@
Linear magnetoelectric and bilinear piezomagnetoelectric NRZSSs — We continue to carry out symmetry analysis and extract the SPGs that host the electric-field switchable NRZSSs, following the logics mentioned in the previous section. Our detailed analysis can be found in Sections I and II of the Supplementary Material (SM), where the SPGs hosting magnetoelectric and piezomagnetoelectric NRZSSs are summarized in Tables S6 and S9 therein. As shown in Eqs. (4) and (6), the $E_{\alpha}$-driven magnetoelectric or piezomagnetoelectric NRZSSs may be rooted in $E_{\alpha}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, $E_{\alpha}^{2}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}^{2}\sigma_{\chi}$, or other higher-order couplings (see Tables S11 and S12 of the SM). Here, we focus on the predominant electric-field induced NRZSSs which include linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs. In Tables 1 and 2, we collect the nonpolar collinear SPGs that enable these two types of NRZSSs.
The linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are characterized by $\lambda_{1}E_{\alpha}\sigma_{\chi}$ and $\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, respectively. At the Brillouin zone center, the electronic eigen energies are $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$, with $\varepsilon_{0}$ being an energy independent of spin. When changing the electric field from $E_{\alpha}$ to $-E_{\alpha}$, the eigen energies become $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$. This means that both linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are electric-field switchable.
Table 1: The nonpolar collinear SPGs that enable linear magnetoelectric NRZSSs. For an SPG associated with the $E_{\alpha}\sigma_{\chi}$ coupling, a $\circ$” symbol is filled in the corresponding entry; Such an entry is endowed with a “--” symbol if the $E_{\alpha}\sigma_{\chi}$ coupling is not symmetrically allowed. For short, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its ${}^{\infty m}1$ part. As an example, the ${}^{1}m^{1}m^{\bar{1}}m^{\infty m}1$ SPG (abbreviated as ${}^{1}m^{1}m^{\bar{1}}m$) hosts the $E_{z}$-driven linear magnetoelectric NRZSSs via the $E_{z}\sigma_{\chi}$ coupling, but does not accommodate the $E_{x}$- or $E_{y}$-driven linear magnetoelectric NRZSSs.
@@
${}^{\bar{1}}\bar{1}$circ $\circ$ $\circ$ ${}^{\bar{1}}2/^{1}m$circ -- $\circ$
${}^{1}2/^{\bar{1}}m$- $\circ$ -- ${}^{\bar{1}}2^{\bar{1}}2^{1}2$- -- $\circ$
${}^{\bar{1}}2^{1}2^{\bar{1}}2$- $\circ$ -- ${}^{1}2^{\bar{1}}2^{\bar{1}}2$circ -- --
${}^{\bar{1}}m^{1}m^{1}m$circ -- -- ${}^{1}m^{\bar{1}}m^{1}m$- $\circ$ --
${}^{1}m^{1}m^{\bar{1}}m$- -- $\circ$ ${}^{\bar{1}}\bar{4}$- -- $\circ$
${}^{1}4/^{\bar{1}}m$- -- $\circ$ ${}^{1}4^{\bar{1}}2^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{4}^{\bar{1}}2^{1}m$- -- $\circ$ ${}^{1}4/^{\bar{1}}m^{1}m^{1}m$- -- $\circ$
${}^{\bar{1}}\bar{3}$- -- $\circ$ ${}^{1}3^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{3}^{1}m$- -- $\circ$ ${}^{\bar{1}}\bar{6}$- -- $\circ$
${}^{1}6/^{\bar{1}}m$- -- $\circ$ ${}^{1}6^{\bar{1}}2^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2$- -- $\circ$ ${}^{1}6/^{\bar{1}}m^{1}m^{1}m$- -- $\circ$
@@
Table 2: The nonpolar collinear SPGs that enable the bilinear piezomagnetoelectric NRZSSs. The $\eta_{xx}$, $\eta_{yy}$, $\eta_{zz}$, $\eta_{xy}$, $\eta_{yz}$, and $\eta_{zx}$ strains are represented by “1”, “2”, “3”, “4”, “5”, and “6” numbers, respectively. For an SPG associated with the $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ coupling, the number representing the $\eta_{\beta\gamma}$ strain is filled in the corresponding entry. If an SPG enables $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ and $E_{\alpha}\eta_{\beta^{\prime}\gamma^{\prime}}\sigma_{\chi}$ couplings, two numbers representing the $\eta_{\beta\gamma}$ and $\eta_{\beta^{\prime}\gamma^{\prime}}$ strains are filled in the corresponding entry, and so forth. The entry with “--” indicates that the corresponding bilinear piezomagnetoelectric NRZSSs are symmetrically forbidden. In this table, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its ${}^{\infty m}1$ part. As an example, the ${}^{\bar{1}}\bar{3}^{1}m^{\infty m}1$ SPG (abbreviated as ${}^{\bar{1}}\bar{3}^{1}m$ in this table) hosts the $E_{x}$-driven bilinear piezomagnetoelectric NRZSSs via $E_{x}\eta_{xy}\sigma_{\chi}$ or $E_{x}\eta_{zx}\sigma_{\chi}$ couplings, but does not accommodate the $E_{y}$- and $E_{z}$-driven bilinear piezomagnetoelectric NRZSSs.
@@
${}^{\bar{1}}2^{\bar{1}}2^{1}2$ $6$ $5$- ${}^{\bar{1}}2^{1}2^{\bar{1}}2$ $4$- $5$
${}^{1}2^{\bar{1}}2^{\bar{1}}2$- $4$ $6$ ${}^{\bar{1}}m^{1}m^{1}m$- $4$ $6$
${}^{1}m^{\bar{1}}m^{1}m$ $4$- $5$ ${}^{1}m^{1}m^{\bar{1}}m$ $6$ $5$-
${}^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$ $4$ ${}^{1}4/^{\bar{1}}m$ $56$ $56$-
${}^{\bar{1}}4/^{\bar{1}}m$ $56$ $56$ $124$ ${}^{\bar{1}}4^{1}2^{\bar{1}}2$- -- $4$
${}^{\bar{1}}4^{\bar{1}}2^{1}2$ $6$ $5$ $12$ ${}^{1}4^{\bar{1}}2^{\bar{1}}2$ $6$ $5$-
${}^{\bar{1}}4^{\bar{1}}2^{1}m$ $6$ $5$- ${}^{1}4^{\bar{1}}2^{\bar{1}}m$ $6$ $5$ $12$
${}^{1}4/^{\bar{1}}m^{1}m^{1}m$ $6$ $5$- ${}^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m$ $5$ $6$ $4$
${}^{\bar{1}}4/^{\bar{1}}m^{1}m^{\bar{1}}m$ $6$ $5$ $12$ ${}^{1}4/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$-
${}^{\bar{1}}\bar{3}^{1}m$ $46$- -- ${}^{\bar{1}}\bar{3}^{\bar{1}}m$- $46$-
${}^{1}6/^{\bar{1}}m$ $56$ $56$- ${}^{\bar{1}}6^{\bar{1}}2^{1}2$ $4$- --
${}^{\bar{1}}6^{1}2^{\bar{1}}2$- $4$- ${}^{1}6^{\bar{1}}2^{\bar{1}}2$ $6$ $5$-
${}^{\bar{1}}\bar{6}^{\bar{1}}m^{1}2$ $5$- -- ${}^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2$ $6$- --
${}^{1}\bar{6}/^{\bar{1}}m^{1}m^{1}m$ $6$ $5$- ${}^{\bar{1}}\bar{6}/^{1}m^{\bar{1}}m^{1}m$- $4$-
${}^{\bar{1}}\bar{6}/^{1}m^{1}m^{\bar{1}}m$ $4$- -- ${}^{1}\bar{6}/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$-
${}^{\bar{1}}m^{\bar{1}}\bar{3}$ $5$ $6$ $4$ ${}^{\bar{1}}4^{1}3^{\bar{1}}2$ $5$ $6$ $4$
${}^{\bar{1}}m^{\bar{1}}\bar{3}^{1}m$ $5$ $6$ $4$

### v2

@@ Abstract
Magnetic or electrical manipulation of electronic spin is elementary for spin-based logic, computing, and memory, where the latter is a low-power manipulation scheme. Rashba-like spin splittings stemming from spin-orbit interaction (SOI) enable electric-field manipulation of spin, but the relativistic SOI causes spin relaxations and yields dissipative transport of spin-encoded information. Recent works suggest the occurrence of electric-field switchable non-relativistic Zeeman spin splittings (NRZSSs) in collinear antiferromagnets — allowing for electrical manipulation of spin in the non-relativistic regime; yet, a theory elucidating the mechanisms for these NRZSSs and guiding the materials discovery remains missing. Here, we develop such a theory by analyzing the symmetries of spin point groups. We highlight the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms for NRZSSs that depend linearly on electric field and are electrically switchable. First-principles calculations further confirm that FeOOH and NaMnP showcase such NRZSSs. Our theory provides guidelines for discovering light-element collinear antiferromagnets with electrically switchable NRZSSs, which are promising for the design of high-performance spin-based devices.
@@
Introduction — Spintronic devices (e.g., logic, computing, and memory) are built upon the manipulation of electronic spin in solids [1, 2, 3, 4, 5]. Such a manipulation can be achieved by magnetic field or electric field, where the latter offers fascinating low-power devices by reducing the “Joule heating” [6, 7, 8]. In principle, electric-field manipulation of spin is enabled by electrically switchable spin splittings [9, 10]. Because of the relativistic spin-orbit interaction (SOI), dielectrics and ferroelectrics naturally host electrically switchable Rashba-like spin splittings [11, 12, 13, 14, 15]. As a by-product, SOI causes the so-called D’yakonov-Perel’ spin relaxation as well [16, 17, 18], which yields the spin decoherence of electrons and prevents the long-range transport of the spin-encoded information [19, 18, 20, 21]. To suppress the D’yakonov-Perel’ spin relaxation, the proposal of electrically switchable spin splittings in the non-relativistic regime is of high necessity.
Recent works highlight two types of notable spin splittings: alternating spin splittings [22, 23, 24, 25, 26, 27, 28, 29] and non-relativistic Zeeman spin splittings (NRZSSs) [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], both of which provide platforms for achieving electrical manipulation of spin within the non-relativistic regime. Alternating spin splittings occur in altermagnets, which can exhibit $d$-, $g$-, or $i$-wave anisotropy in the Brillouin zone [22, 23, 24, 25], and in certain cases, such splittings can be controlled by electric fields [27, 28, 29]. Unlike the alternating spin splittings, NRZSSs lift spin degeneracy uniformly across the entire Brillouin zone [35, 43]. This feature enables the generation of fully spin-polarized currents [45, 35, 43], which is ideal for highly efficient spin injection in spintronic devices [46, 47]. Notably, theoretical works suggest the possibility of achieving NRZSSs in collinear antiferromagnets that are switchable by an electric field [36, 37, 38, 43, 39, 40, 41, 42, 44]. For example, electrically switchable NRZSSs were predicted in two-dimensional layer materials (e.g., bilayer FeBr2 [42] and bilayer VSe2 [36]) and bulk materials (e.g., Fe2TeO6 [37]). Regarding electrically switchable Zeeman spin splittings, a theory was developed by analyzing the magnetic point groups (involving the SOI) [37]. Such a theory, rooted in the relativistic regime, can neither interpret the aforementioned NRZSSs nor predict materials with electrically switchable NRZSSs.
@@
So far, mechanisms for electric-field switchable NRZSSs are elusive and selection rules for the corresponding materials discovery are missing. To address this issue, we analyze the symmetries of spin point groups (SPGs) that are associated with collinear magnets in the non-relativistic limit. This allows us to develop a theory on electrically switchable NRZSSs in collinear antiferromagnets, and reveal the magnetoelectric and piezomagnetoelectric mechanisms for this phenomenon. Our theory provides theoretical guidelines for screening collinear antiferromagnets with electrically switchable NRZSSs. This yields the first-principles predictions that FeOOH and NaMnP are two candidates with linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs, respectively.
NRZSSs in collinear magnets — Without the inclusion of SOI, the electronic spin angular momentum is conserved in collinear magnets  [^1] To design spin-based devices, spin conservation is a desired feature [19]. While spin conservation is a merit of non-relativistic collinear magnets, the occurrence of spin conservation is not guaranteed in non-relativistic noncollinear magnets (see e.g., Refs. [49, 50, 23]). The non-relativistic noncollinear magnets are thus not of interest to us. and serves as a good quantum number to label the electronic eigen states therein [49, 50, 23, 24]. In this case, each spin- and momentum-resolved eigen state is associated with an eigen energy of $\varepsilon(\mathbf{k},s_{\chi})$, where $\mathbf{k}$ is the electronic momentum, $s_{\chi}$ the spin angular momentum, and $\chi$ the characteristic direction for collinear magnets (i.e., magnetic moments being along $\bm{\chi}$ or $\bm{\chi}$). According to Refs. [51, 52, 53, 54], the symmetries of non-relativistic collinear magnets can be well described by 32 ferromagnetic or ferrimagnetic SPGs and 58 antiferromagnetic SPGs. The symmetry operations in SPGs are represented by $R_{s}\left|\right|R_{l}]$, with the spin operation $R_{s}$ and spatial operation $R_{l}$ being separated by the $\left|\right|$” symbol [51, 52]. The $R_{s}\left|\right|R_{l}]$ symmetry operation links the $(\mathbf{k},s_{\chi})$ and $(R_{l}R_{s}\mathbf{k},R_{s}s_{\chi})$ eigen states  [^2] The spin operation can be written as $R_{s}=U_{s}$ or $R_{s}=U_{s}T$, with $U_{s}$ and $T$ being spin proper rotation and time reversal operation, respectively. According to Refs. [51, 52], $U_{s}\mathbf{k}=\mathbf{k}$ and $U_{s}T\mathbf{k}=-\mathbf{k}$., implying the $\varepsilon(\mathbf{k},s_{\chi})=\varepsilon(R_{l}R_{s}\mathbf{k},R_{s}s_{\chi})$ relation. At the Brillouin zone center, both $\mathbf{k}=0$ and $R_{l}R_{s}\mathbf{k}=0$ are valid so that the aforementioned relation can be simplified as $\varepsilon(s_{\chi})=\varepsilon(R_{s}s_{\chi})$ by omitting $\mathbf{k}$.
@@
For all the $R_{s}\left|\right|R_{l}]$ symmetry operations in ferromagnetic or ferrimagnetic SPGs, $R_{s}s_{\chi}=s_{\chi}$ is always true and $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ is not necessarily enforced [51, 52]. This means that non-relativistic collinear ferromagnets or ferrimagnets enable the spontaneous NRZSSs (see e.g., Refs. [50, 23, 24]). Unlike those in ferromagnetic or ferrimagnetic SPGs, the $R_{s}\left|\right|R_{l}]$ symmetry operations in an antiferromagnetic SPG can be partitioned into two subsets [51, 52, 23, 24], namely,
@@
and the $\alpha$” superscript labels the symmetry reduction caused by the $E_{\alpha}$ electric field  [^3] When $E_{\alpha}$ is not compatible with the spatial $R_{l}$ operation (i.e., $R_{l}E_{\alpha}\neq E_{\alpha}$), the application of $E_{\alpha}$ break the $R_{s}\left|\right|R_{l}]$ symmetry operation.. When $L_{-}^{\alpha}$ becomes empty, there will be no symmetry operations that enforce the $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ degeneracy — enabling the electric-field induced NRZSSs. The $E_{\alpha}$-field-induced NRZSSs are described by the effective Hamiltonian
@@
Table 1: The nonpolar collinear SPGs that enable linear magnetoelectric NRZSSs. For an SPG associated with the $E_{\alpha}\sigma_{\chi}$ coupling, a $\circ$” symbol is filled in the corresponding entry; Such an entry is endowed with a “-” symbol if the $E_{\alpha}\sigma_{\chi}$ coupling is not symmetrically allowed. For short, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its $^{\infty m}1$ part. As an example, the $^{1}m^{1}m^{\bar{1}}m^{\infty m}1$ SPG (abbreviated as $^{1}m^{1}m^{\bar{1}}m$) hosts the $E_{z}$-driven linear magnetoelectric NRZSSs via the $E_{z}\sigma_{\chi}$ coupling, but does not accommodate the $E_{x}$- or $E_{y}$-driven linear magnetoelectric NRZSSs.
@@
$^{\bar{1}}\bar{1} \circ \circ \circ$ $^{\bar{1}}2/^{1}m \circ - \circ$
$^{1}2/^{\bar{1}}m - \circ -$ $^{\bar{1}}2^{\bar{1}}2^{1}2 - - \circ$
$^{\bar{1}}2^{1}2^{\bar{1}}2 - \circ -$ $^{1}2^{\bar{1}}2^{\bar{1}}2 \circ - -$
$^{\bar{1}}m^{1}m^{1}m \circ - -$ $^{1}m^{\bar{1}}m^{1}m - \circ -$
$^{1}m^{1}m^{\bar{1}}m - - \circ$ $^{\bar{1}}\bar{4} - - \circ$
$^{1}4/^{\bar{1}}m - - \circ$ $^{1}4^{\bar{1}}2^{\bar{1}}2 - - \circ$
$^{\bar{1}}\bar{4}^{\bar{1}}2^{1}m - - \circ$ $^{1}4/^{\bar{1}}m^{1}m^{1}m - - \circ$
$^{\bar{1}}\bar{3} - - \circ$ $^{1}3^{\bar{1}}2 - - \circ$
$^{\bar{1}}\bar{3}^{1}m - - \circ$ $^{\bar{1}}\bar{6} - - \circ$
$^{1}6/^{\bar{1}}m - - \circ$ $^{1}6^{\bar{1}}2^{\bar{1}}2 - - \circ$
$^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2 - - \circ$ $^{1}6/^{\bar{1}}m^{1}m^{1}m - - \circ$
Linear magnetoelectric and bilinear piezomagnetoelectric NRZSSs — We continue to carry out symmetry analysis and extract the SPGs that host the electric-field switchable NRZSSs, following the logics mentioned in the previous section. Our detailed analysis can be found in Sections I and II of the Supplementary Material (SM)  [^5] See Supplemental Material at [URL] for the analysis of spin point groups, the analysis of electric-field induced non-relativistic Zeeman spin splittings, methods, some numerical results, and the analysis of FeOOH., which includes Refs. [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. The SPGs hosting magnetoelectric and piezomagnetoelectric NRZSSs are summarized in Tables S6 and S9 of the SM. As shown in Eqs. (4) and (6), the $E_{\alpha}$-driven magnetoelectric or piezomagnetoelectric NRZSSs may be rooted in $E_{\alpha}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, $E_{\alpha}^{2}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}^{2}\sigma_{\chi}$, or other higher-order couplings (see Tables S11 and S12 of the SM). Here, we focus on the predominant electric-field induced NRZSSs which include linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs. In Tables 1 and 2, we collect the nonpolar collinear SPGs that enable these two types of NRZSSs.
The linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are characterized by $\lambda_{1}E_{\alpha}\sigma_{\chi}$ and $\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, respectively. At the Brillouin zone center, the electronic eigen energies are $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$, with $\varepsilon_{0}$ being an energy independent of spin. When changing the electric field from $E_{\alpha}$ to $E_{\alpha}$, the eigen energies become $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$. This means that both linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are electric-field switchable.
@@
Table 2: The nonpolar collinear SPGs that enable the bilinear piezomagnetoelectric NRZSSs. The $\eta_{xx}$, $\eta_{yy}$, $\eta_{zz}$, $\eta_{xy}$, $\eta_{yz}$, and $\eta_{zx}$ strains are represented by “1”, “2”, “3”, “4”, “5”, and “6” numbers, respectively. For an SPG associated with the $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ coupling, the number representing the $\eta_{\beta\gamma}$ strain is filled in the corresponding entry. If an SPG enables $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ and $E_{\alpha}\eta_{\beta^{\prime}\gamma^{\prime}}\sigma_{\chi}$ couplings, two numbers representing the $\eta_{\beta\gamma}$ and $\eta_{\beta^{\prime}\gamma^{\prime}}$ strains are filled in the corresponding entry, and so forth. The entry with “-” indicates that the corresponding bilinear piezomagnetoelectric NRZSSs are symmetrically forbidden. In this table, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its $^{\infty m}1$ part. As an example, the $^{\bar{1}}\bar{3}^{1}m^{\infty m}1$ SPG (abbreviated as $^{\bar{1}}\bar{3}^{1}m$ in this table) hosts the $E_{x}$-driven bilinear piezomagnetoelectric NRZSSs via $E_{x}\eta_{xy}\sigma_{\chi}$ or $E_{x}\eta_{zx}\sigma_{\chi}$ couplings, but does not accommodate the $E_{y}$- and $E_{z}$-driven bilinear piezomagnetoelectric NRZSSs.
@@
$^{\bar{1}}2^{\bar{1}}2^{1}2$ 66 55 - $^{\bar{1}}2^{1}2^{\bar{1}}2$ 44 - 55
$^{1}2^{\bar{1}}2^{\bar{1}}2 -$ 44 66 $^{\bar{1}}m^{1}m^{1}m -$ 44 66
$^{1}m^{\bar{1}}m^{1}m$ 44 - 55 $^{1}m^{1}m^{\bar{1}}m$ 66 55 -
$^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ 55 66 44 $^{1}4/^{\bar{1}}m$ $56$ $56$ -
$^{\bar{1}}4/^{\bar{1}}m$ $56$ $56$ $124$ $^{\bar{1}}4^{1}2^{\bar{1}}2 - -$ 44
$^{\bar{1}}4^{\bar{1}}2^{1}2$ 66 55 $12$ $^{1}4^{\bar{1}}2^{\bar{1}}2$ 66 55 -
$^{\bar{1}}\bar{4}^{\bar{1}}2^{1}m$ 66 55 - $^{1}\bar{4}^{\bar{1}}2^{\bar{1}}m$ 66 55 $12$
$^{1}4/^{\bar{1}}m^{1}m^{1}m$ 66 55 - $^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m$ 55 66 44
$^{\bar{1}}4/^{\bar{1}}m^{1}m^{\bar{1}}m$ 66 55 $12$ $^{1}4/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ 55 66 -
$^{\bar{1}}\bar{3}^{1}m$ $46$ - - $^{\bar{1}}\bar{3}^{\bar{1}}m -$ $46$ -
$^{1}6/^{\bar{1}}m$ $56$ $56$ - $^{\bar{1}}6^{\bar{1}}2^{1}2$ 44 - -
$^{\bar{1}}6^{1}2^{\bar{1}}2 -$ 44 - $^{1}6^{\bar{1}}2^{\bar{1}}2$ 66 55 -
$^{\bar{1}}\bar{6}^{\bar{1}}m^{1}2$ 55 - - $^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2$ 66 - -
$^{1}6/^{\bar{1}}m^{1}m^{1}m$ 66 55 - $^{\bar{1}}6/^{1}m^{\bar{1}}m^{1}m -$ 44 -
$^{\bar{1}}6/^{1}m^{1}m^{\bar{1}}m$ 44 - - $^{1}6/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ 55 66 -
$^{\bar{1}}m^{\bar{1}}\bar{3}$ 55 66 44 $^{\bar{1}}4^{1}3^{\bar{1}}2$ 55 66 44
$^{\bar{1}}m^{\bar{1}}\bar{3}^{1}m$ 55 66 44
Materials with electric-field switchable NRZSSs — We recall that the electrically switchable NRZSSs in bilayer FeBr2, bilayer VSe2, bulk Fe2TeO6 and other materials, previously predicted in Refs. [36, 37, 38, 43, 39, 40, 41, 42, 44], can be ascribed to the linear magnetoelectric mechanisms in our theory (see Section II.D of the SM). Next, we explore more materials that might showcase electric-field switchable NRZSSs (i.e., linear magnetoelectric or bilinear piezomagnetoelectric NRZSSs). This is assisted by the MAGNDATA database [77, 78], and guided by Tables 1 and 2. We search for nonpolar collinear antiferromagnets composed of light elements, identify the SPGs for such antiferromagnets, and discover the FeOOH and NaMnP as two representatives with sizeable electric-field switchable NRZSSs. [The strength of SOI basically depends on the atomic mass. Materials composed of heavy elements (e.g., Bi, Sb, and Pt) likely showcase sizable SOI [80, 81, 82].] Especially, both FeOOH and NaMnP are room-temperature antiferromagnets, with Néel temperature being >>343 K [83, 84] for FeOOH and >>293 K [85, 86] for NaMnP.